OpenStudy (anonymous):

two similar cones have heights 10 cm and 15 cm. Find the ratios of each of the following. a. The radii b. the surface areas of the cones c. the volume of the cones

OpenStudy (anonymous):

All 'lengths' (i.e. height, radii) have the same ratio, let's call it k. All 'areas' have the ratio k^2. The volume has the ratio k^3.

OpenStudy (anonymous):

I dont get it

OpenStudy (anonymous):

15/10 = 3/2 or 3:2 ratio. Let this be 'k', and the results follow from above.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

how do you find the radi when it only gives me the heigt

OpenStudy (anonymous):

You don't need to find the radii, only their ratio. As they are similar, the ratio of the radii is equal to the ratio of the heights.

OpenStudy (anonymous):

so the radii would be 3/2

OpenStudy (anonymous):

The RATIO of the radii is 3/2 or 3:2, yes.

OpenStudy (anonymous):

When I have to do the equation its pie r s + pie r2 what do i subsstitue fro r

OpenStudy (anonymous):

You can't do that. And you don't need to. It is really as simple as my first reply.

OpenStudy (anonymous):

My teacher didnt teach me that way, to me that method is coufusing idk how to find the surface area and volume

OpenStudy (anonymous):

You CANNOT find them explicitly. The description of 'height 10' could be any cone with any radius.

OpenStudy (anonymous):

so how do i find the s.a and volume

OpenStudy (anonymous):

The ratio of the SA = (3/2)^2 = 9/4 The ratio of the Volume = (3/2)^2 = 27/8 You CANNOT find the exact SA/volume.

OpenStudy (anonymous):

K THNX can you helop me on my othere question

OpenStudy (anonymous):

two similar pyramids have volumes 3 in3 and 375 in3. find the ratios of each of the following. a. the heights of the pyramids b the lateral areas of the pyramids c. the surface areas of the pyramids